OpenStudy (anonymous):

Simplify 27^2/3 by hand (without using a calculator)

OpenStudy (anonymous):

\[27^{2/3}=(27^{1/3})^{2}=3^2=9\]

OpenStudy (anonymous):

That's exactly how I would have done it.